4. A corporation had 10,000 shares of $10 par value common stock outstanding when the board of directors declared a stock dividend of 30%. At the time of the stock dividend, the market value per share was $12. The entry to record this dividend is: A. Debit Retained Earnings $36,000; credit Common Stock $36,000. B. Debit Retained Earnings $36,000; credit Common Stock $30,000, credit Paid-In Capital in Excess of Par Value $6,000. C. Debit Common Stock $36,000; credit Retained Earnings $36,000. D. Debit Retained Earnings $30,000; credit Common Stock $30,000. E. No entry is needed.
Answer #1

Stock dividend or decrease in retained earnings =10,000×$12×30% = $36,000

Common stock = 10,000×$10×30% = $30,000

Paid in capital in excess of par = 36,000-30,000 = $6,000

Debit retained earnings $36,000; credit common stock $30,000; credit paid in capital in excess of par value $6,000.

Hence, the correct answer is option b.